This came from Factoring Polynomial Section
Solve
x^2+(x+7)^2=17^2

Answer by stanbon(75887)

x^2+(x+7)^2=17^2
x^2 + x^2 + 14x + 49 = 289
2x^2 + 14x - 240 = 0
x^2 + 7x - 120 = 0
x^2 + 15x - 8x - 120 = 0
x(x+15) - 8(x + 15) = 0
(x+15)(x-8) = 0
x = -15 or x = 8
====================
Cheers,
Stan H.